# From Lab to Field: Real-World Evaluation of an AI-Driven Smart Video Solution to Enhance Community Safety

### v2

@@ Abstract
This article adopts and evaluates an AI-enabled Smart Video Solution (SVS) designed to enhance safety in the real world. The system integrates with existing infrastructure camera networks, leveraging recent advancements in AI for easy adoption. Prioritizing privacy and ethical standards, pose based data is used for downstream AI tasks such as anomaly detection. Cloud-based infrastructure and mobile app are deployed, enabling real-time alerts within communities. The SVS employs innovative data representation and visualization techniques, such as the Occupancy Indicator, Statistical Anomaly Detection, Bird’s Eye View, and Heatmaps, to understand pedestrian behaviors and enhance public safety. Evaluation of the SVS demonstrates its capacity to convert complex computer vision outputs into actionable insights for stakeholders, community partners, law enforcement, urban planners, and social scientists. This article presents a comprehensive real-world deployment and evaluation of the SVS, implemented in a community college environment across 16 cameras. The system integrates AI-driven visual processing, supported by statistical analysis, database management, cloud communication, and user notifications. Additionally, the article evaluates the end-to-end latency from the moment an AI algorithm detects anomalous behavior in real-time at the camera level to the time stakeholders receive a notification. The results demonstrate the system’s robustness, effectively managing 16 CCTV cameras with a consistent throughput of 16.5 frames per second (FPS) over a 21-hour period and an average end-to-end latency of 26.76 seconds between anomaly detection and alert issuance.
Introduction
Figure 1: The end-to-end system architecture coupled with notification’s data flow. $S_{1}$ represents the first scene that detects a behavior anomaly. The yellow line shows the notification data flow to the notification service. $S_{n}$ shows the $n^{\text{th}}$ scene where a suspicious object has been detected. The red line shows the notification data flow when detecting object anomalies.
@@ Introduction
Closed-Circuit Television (CCTV) cameras are widely applied in public spaces, traditionally serving as passive surveillance tools aimed primarily at security and monitoring (Ashby 2017). However, these systems often underutilized the potential of visual data they capture, missing opportunities to contribute to broader civic applications such as urban planning, public health, and resource management (Angelidou et al. 2018). By transforming CCTV networks into Smart Video Solutions (SVS), we can unlock new functionalities that go beyond mere surveillance to enhance community living and environments (Javed et al. 2022).
@@
SVS leverage advancements in AI and machine learning to analyze video data in real-time, providing actionable insights that can improve city infrastructure and community services (Ardabili et al. 2023b). For example, the results can be used to inform urban planners about pedestrian flows for better traffic management (Rashvand et al. 2023), help public health officials monitor social distancing (Usmani et al. 2023), and assist in the efficient allocation of community resources (Dechouniotis et al. 2020). Such applications demonstrate the potential of SVS to transform video data into valuable information for civic improvements.
@@
While traditional research has focused on the development of these systems in controlled environments (Pazho et al. 2023a), it is crucial to transition and evaluate SVS in real-world settings to address challenges such as latency, scalability, and privacy concerns (Ardabili et al. 2023a; Xu et al. 2018). Moving from lab to field ensures that these systems are robust and adaptable, capable of handling the complexities of dynamic urban environments and diverse community needs (Noghre et al. 2023).
@@
The architecture of the SVS is illustrated in Figure 1, which represents the end-to-end design (Ardabili et al. 2023b). It begins with the environment where video data is captured by existing CCTV cameras. The data is then processed by the AI module, where computer vision techniques are applied for human behavior analysis. The processed information is sent to the server module for further data analysis and aggregation (Pazho et al. 2023a). To ensure privacy, metadata is stored in the cloud module, which also hosts a mobile application. This setup allows end-users to receive notifications and access insights on their devices, fostering real-time engagement and decision-making.
@@
This article introduces the deployment and evaluation of an AI-enabled SVS designed to enhance safety and civic functionalities in community spaces such as educational areas, parking lots, and smart city applications. By employing privacy-preserving techniques, the system utilizes pose based data for human behavior analysis, aligning with ethical standards. The SVS applies innovative data visualization tools, including Occupancy Indicators, Statistical Anomaly Detection, Bird’s Eye View, and Heatmaps, to provide insights into pedestrian behaviors and resource allocation, thereby aiding urban planning and enhancing public safety.
@@
  • Adopting an AI-enabled SVS with existing CCTV infrastructure in real world, emphasizing civic applications beyond traditional security measures.
@@
Related Works
@@
Recent advancements in AI-driven video systems have leveraged state-of-the-art algorithms and system designs to perform high-level AI tasks using computer vision. Ancilia, for instance, employs a multi-stage computer vision pipeline aimed at enhancing public safety. While we have implemented this system as the baseline for our deployment, it is important to note that the original study reports the results in controlled environment. (Pazho et al. 2023a).
@@
The deployment of SVS in real-world settings is essential to validate their effectiveness and address challenges such as latency, scalability, and privacy. While traditional research has focused on laboratory-based studies for real-time object tracking and anomaly detection (Pazho et al. 2023b; S. et al. 2021), there is a growing recognition of the need to evaluate these systems in practical environments. Real-world evaluation ensures that this technology can adapt to dynamic urban settings and meet diverse community needs.
@@
Recent advancements in machine learning have enabled smart video technology to perform complex data analysis, offering insights for civic applications. For instance, Pramanik et al. demonstrated lightweight models for edge devices to reduce latency in urban environments (Pramanik, Sarkar, and Maiti 2021). Singh et al. introduced an AI-based video technology for fall detection, highlighting the potential for applications beyond traditional safety measures (Singh et al. 2023).
@@
In 2023, the E2E-VSDL method by Gandapur utilized BiGRU and CNN for anomaly detection, achieving high accuracy (Gandapur 2022). Similarly, TAC-Net’s deep learning approach excelled in addressing anomaly scenarios (Huang et al. 2021).
@@
Integrating these systems into smart city applications bridges the gap between theoretical research and practical implementation. Alshammari et al. developed an SVS using surveillance cameras in real-world settings (Alshammari and Rawat 2019), and RV College of Engineering Bengaluru improved object detector accuracy (Franklin and Dabbagol 2020). These efforts highlight the necessity of robust testbed support for evaluating SVS, addressing scalability and privacy challenges (Ma et al. 2019).
@@
The combination of machine learning and IoT devices has transformed urban planning, providing insights for managing congestion and enhancing civic participation. Ardabili et al. introduced an end-to-end system design that integrates Ancilia with an end-user device for direct communication. We have adopted this comprehensive system design for our study’s deployment, allowing us to assess its effectiveness in practical, real-world environments (Ardabili et al. 2023b). Mahdavinejad et al. emphasized machine learning’s role in forecasting urban congestion (Mahdavinejad et al. 2018), while Zanella et al. advocated for accessible urban IoT data (Zanella et al. 2014). These studies illustrate how data-driven approaches contribute to civic applications.
@@
Figure 2: End-to-end detailed system. $C_{0}$ represents the camera ID and for each camera, one AI Module $N_{0}$ including multi-AI-vision-models pipeline is assigned with. All the AI Modules send processed data to one Server Module database, and Server Module will re-identify the human track ID based on the feature extractor data. The statistical analyzer analyzes all the data in the database stored across all the cameras and communicates the results with Cloud Module. Cloud-native services are utilized to host the end users’ applications.
Despite progress in SVS research, comprehensive testbed support for evaluating real-world system performance remains limited. Our study addresses this gap by deploying a state-of-the-art SVS in a community college, demonstrating its capabilities and potential for enhancing civic applications in public settings.
Software System Features
The AI-based real-time video solution seamlessly integrates with existing CCTV infrastructures, creating a Physical-Cyber-Physical (PCP) system that delivers actionable information to end users. Such system consists of four key components: AI Modules, Server Module, Cloud Module, and End-user Devices, as illustrated in Figure 2. The design of the AI and server modules was inspired by work (Pazho et al. 2023a), while the cloud module and user device components were adapted from another study (Ardabili et al. 2023b).
AI Module/Modules
@@
is a multi-stage pipeline optimized for real-time computer vision tasks. It processes image data in batches of 30 frames, where an object detector identifies objects and flags anomalies (Pazho et al. 2023a). A tracking algorithm then creates tracklets, and pose estimator extracts 2D skeletal data for individuals. Object anomaly detection and behavioral anomaly detection (Noghre, Pazho, and Tabkhi 2024; Pazho et al. 2024) are performed within this module respectively, with alerts sent to end-user devices for real-time response (Pazho et al. 2023b; Yao et al. 2024).
In our feature, YOLOv8 (Jocher, Chaurasia, and Qiu 2023), Bytetrack (Zhang et al. 2022), HRNet (Sun et al. 2019), GEPC (Markovitz et al. 2020a) and OSNet (Zhou et al. 2019) are used for further evaluation. The detailed decision-making process is provided in the supplementary materials for replication and further investigation.
Server Module
centralizes metadata and historical data from AI modules, stored in a MySQL database as outlined in Table 1. It handles global tracking and statistical analysis, using cosine similarity to re-identify individuals across cameras to analyze patterns (Pazho et al. 2023a). The design prioritizes privacy by avoiding raw data transmission to the Cloud Module.

### v3

@@ Abstract
This article adopts and evaluates an AI-enabled Smart Video Solution (SVS) designed to enhance safety in the real world. The system integrates with existing infrastructure camera networks, leveraging recent advancements in AI for easy adoption. Prioritizing privacy and ethical standards, pose-based data is used for downstream AI tasks such as anomaly detection. A Cloud-based infrastructure and a mobile app are deployed, enabling real-time alerts within communities. The SVS employs innovative data representation and visualization techniques, such as the Occupancy Indicator, Statistical Anomaly Detection, Bird’s Eye View, and Heatmaps, to understand pedestrian behaviors and enhance public safety. Evaluation of the SVS demonstrates its capacity to convert complex computer vision outputs into actionable insights for stakeholders, community partners, law enforcement, urban planners, and social scientists. This article presents a comprehensive real-world deployment and evaluation of the SVS, implemented in a community college environment with 16 cameras. The system integrates AI-driven visual processing, supported by statistical analysis, database management, cloud communication, and user notifications. Additionally, the article evaluates the end-to-end latency from the moment an AI algorithm detects anomalous behavior in real-time at the camera level to the time stakeholders receive a notification. The results demonstrate the system’s robustness, effectively managing 16 CCTV cameras with a consistent throughput of 16.5 frames per second (FPS) over a 21-hour period and an average end-to-end latency of 26.76 seconds between anomaly detection and alert issuance.
keywords:
Smart Video Solution, Computer Vision, Case Study, Public Safety, Anomaly Detection
journal: Internet of Things
[…]
[label1]organization=Department of Electrical and Computer Engineering, University of North Carolina at Charlotte,city=Charlotte, state=NC, country=USA
[…]
[label2]organization=Department of Public Policy, University of North Carolina at Charlotte,city=Charlotte, state=NC, country=USA
1 Introduction
@@
Closed-Circuit Television (CCTV) cameras are widely applied in public spaces, traditionally serving as passive surveillance tools aimed primarily at security and monitoring [1]. However, these systems often underutilizing the potential of visual data they capture, missing opportunities to contribute to broader civic applications such as urban planning, public health, and resource management [2]. By transforming CCTV networks into Smart Video Solutions (SVS), we can unlock new functionalities that go beyond mere surveillance to enhance community living and environments [3].
@@
SVS utilizes advanced AI and machine learning techniques to conduct real-time analysis of video data, thereby generating actionable insights that enhance urban infrastructure and community services [4]. Specifically, these insights can inform urban planners about pedestrian flows to improve traffic management [5], enable public health officials to monitor social distancing [6], and facilitate the efficient allocation of community resources [7]. Collectively, these applications underscore the significant potential of SVS to transform video data into valuable information for civic improvements.
@@
While previous research has focused on developing SVS in controlled environments [8], transitioning and evaluating these systems in real-world settings is crucial to address challenges such as latency, scalability, and privacy concerns [9, 10]. This shift from controlled laboratory settings to diverse real-world environments ensures that SVS is robust against real-world variability and adaptable to changing conditions, enabling it to effectively handle the complexities of dynamic urban environments and diverse community needs [11]. This study primarily focuses on evaluating of such a system in real-world settings, leveraging both the AI [8] and application [4] components from the earlier State-Of-The-Art (SOTA) research that emphasized the development of such a system. Through this real-world evaluation, we are assessing the system’s impact in community settings and its broader civic applications.
Figure 1: The end-to-end system architecture is coupled with the notification’s data flow. $S_{1}$ represents the first scene that detects a behavior anomaly. The yellow line shows the notification data flow to the notification service. $S_{n}$ shows the $n^{\text{th}}$ scene where a suspicious object has been detected. The red line shows the notification data flow when detecting object anomalies.
@@
The architecture of the SVS is illustrated in Fig. 1, which represents the end-to-end design [4]. It begins with the environment where video data is captured by existing CCTV cameras. The data is then processed by the AI module, where computer vision techniques are applied for human behavior analysis. The processed information is sent to the server module for further data analysis and aggregation [8]. To ensure privacy, metadata is stored in the cloud module, which also hosts a mobile application. This setup allows end-users to receive notifications and access insights on their devices, fostering real-time engagement and decision-making.
@@
Previous studies have mainly focused on the development of SVS either from the technical or the application perspectives. However, their approaches were mainly limited to lab settings, which lacked real-world validations. Deploying these systems in the real world, introduce new challenges such as network latency, integration with the existing infrastructure, and scalability. Moreover, the main scope of SVS focuses on security applications like detecting behavior anomalies or detecting fire hazards. Generally, there is less debate on privacy in these applications, where the focus is on immediate identification of security threats. However, when there is a need for continuous monitoring of public spaces in broader civic applications, privacy concerns become significantly more important. This further underscores the need to assess real-world performance to ensure that the system operates effectively while complying with ethical aspects. In this study, we bridge the gap in deploying and evaluating an SVS in the real-world setup and go beyond security applications, to show the scalability and robustness in broader civic applications. This article introduces the deployment and evaluation of an AI-enabled SVS designed to enhance safety and civic functionalities in community spaces such as educational areas, parking lots, and smart city applications. By employing privacy-preserving techniques, the system utilizes pose-based data for human behavior analysis, aligning with ethical standards. The SVS applies innovative data visualization tools, including Occupancy Indicators, Statistical Anomaly Detection, Bird’s Eye View, and Heatmaps, to provide insights into pedestrian behaviors and resource allocation, thereby aiding urban planning and enhancing public safety.
@@
  • 1.
    Adopting an AI-enabled SVS with existing CCTV infrastructure in the real world, emphasizing civic applications beyond traditional security measures.
  • 2.
@@
The remainder of this paper is structured as follows. Section 2 provides a comprehensive review of relevant literature in the field. Section 3 outlines the key features and architecture of our system. In Section 4, we detail the implementation methodology and evaluation framework. Section 5 presents our data visualization tools and techniques, while Section 6 examines community engagement and its impact. Section 7 describes our experimental setup and discusses the results. Finally, Section 8 summarizes our findings and discusses future research directions.
2 Related Works
@@
Recent advancements in AI-driven video systems have harnessed state-of-the-art algorithms and system designs to execute high-level AI tasks through computer vision. For example, Ancilia [8] utilizes a multi-stage computer vision pipeline specifically engineered to enhance public safety. Although we adopted this system as the baseline for our deployment, the original study reports results exclusively from a controlled environment. This limitation underscores the necessity of real-world evaluations to confirm the system’s practical efficacy.
@@
Deploying SVS systems in real-world environments is critical to validate their performance and tackle key challenges, including latency, scalability, and privacy concerns. Historically, research has concentrated on laboratory-based studies for real-time object tracking and anomaly detection [12, 13]. However, there is increasing acknowledgment of the need to assess these systems in dynamic, real-world settings. Such evaluations are vital to ensure SVS technologies can adapt to the complexities of urban landscapes and fulfill diverse community requirements.
@@
Recent advancements in machine learning have enabled smart video technology to perform complex data analysis, offering insights for civic applications. For instance, Pramanik et al. demonstrated lightweight models for edge devices to reduce latency in urban environments [14]. Singh et al. introduced an AI-based video technology for fall detection, highlighting the potential for applications beyond traditional safety measures [15].
@@
In 2023, the E2E-VSDL method by Gandapur utilized BiGRU and CNN for anomaly detection, achieving high accuracy [16]. Similarly, TAC-Net’s deep learning approach excelled in addressing anomaly scenarios [17].
@@
Integrating these systems into smart city applications bridges the gap between theoretical research and practical implementation. Alshammari et al. developed an SVS using surveillance cameras in real-world settings [18], and RV College of Engineering Bengaluru improved object detector accuracy [19]. These efforts highlight the necessity of robust testbed support for evaluating SVS, addressing scalability and privacy challenges [20]. Y. Yuan et al. [21] introduced an innovative decentralized framework for Cyber-Physical systems, demonstrating its effectiveness in optimizing traffic flow and reducing pedestrian waiting times in real-world scenarios. While [22] presents an impressive self-aware Cyber-Physical system, it notably lacks comprehensive data on system latency—a crucial metric in testbed environments.
Recent research has increasingly focused on optimizing large-scale Smart Vehicle systems, with significant attention given to enhancing data transformation, communication efficiency, and the integration of blockchain technology across various domains [23]. These developments demonstrate a shift beyond traditional model performance metrics to address comprehensive system scalability and real-world implementation challenges.
@@
The combination of machine learning and IoT devices has transformed urban planning, providing insights for managing congestion and enhancing civic participation. Ardabili et al. introduced an end-to-end system design that integrates Ancilia with an end-user device for direct communication. We have adopted this comprehensive system design for our study’s deployment, allowing us to assess its effectiveness in practical, real-world environments [4]. Mahdavinejad et al. emphasized machine learning’s role in forecasting urban congestion [24], while Zanella et al. advocated for accessible urban IoT data [25]. These studies illustrate how data-driven approaches contribute to civic applications.
Despite significant advances in Smart Vehicle Systems research, there remains a critical gap in the comprehensive testbed evaluation of real-world SVS performance. To address this limitation, we have developed and implemented an SVS system within a public community college setting, providing a practical demonstration of these technologies in a real-world environment. The following sections detail our implementation, findings, and their implications for future SVS deployments.
3 Software System Features
@@
Figure 2: End-to-end detailed system. $C_{0}$ represents the camera ID and for each camera, one AI Module $N_{0}$ including a multi-AI-vision-models pipeline is assigned. All the AI Modules send processed data to one Server Module database, and the Server Module will re-identify the human track ID based on the feature extractor data. The statistical analyzer analyzes all the data stored in the database across all the cameras
and communicates the results with the Cloud Module. Cloud-native services are utilized to host the end users’ applications.
Algorithm 1 SVS Algorithm
1:Initialization:
2:  Initialize AI modules, Server Module, and Cloud Module.
3:for each camera $c_{i}$ in the set of cameras do
4:  while stream from $c_{i}$ is active do
5:   Step 1: Object Detection
6:    Fetch a batch of 30 frames of rgbrgb data from $c_{i}$ as initial input
7:    Apply the object detector to identify humans and objects of interest.
8:    Record bounding boxes $B$ with associated confidence scores.
9:   if objects of interest confidence score in $B\geq\textit{object\_threshold}$ then
10:    Trigger an alert for the detected object of interest to end-user device(s).
11:   else
12:    Continue
13:   end if
14:   Step 2: Tracking
15:    Input $B$ to the tracker to assign unique IDs and generate tracklets $T$.
16:   Step 3: Pose Estimation
17:    For each tracklet in $T$, extract the corresponding input rgbrgb data.
18:    Apply the pose estimator to compute skeletal keypoints $K$.
19:   Step 4: Anomaly Score Prediction
20:    Organize input keypoints into movement sequences.
21:    Compute anomaly scores $\alpha$ for the batch of keypoints.
22:   Step 5: Feature Extraction & Cross-Camera Identification
23:    Extract feature $F$ from input rgbrgb for cross-camera identification.
24:    Transmit $F$ and $\alpha$ to the Server Module for global analyze.
25:   Step 6: Behavior Anomaly Detection
26:    Transmit $\alpha$ to the Server Module for behavior anomaly detection.
27:   if $\alpha\geq\textit{threshold}$ for three consecutive batches then
28:    Trigger an alert to the end-user device(s).
29:   else
30:    Continue
31:   end if
32:   Step 7: Data Processing at Server Module
33:    Process $F$, anomaly scores $\alpha$, and related data for statistical analysis.
34:   Step 8: Data Transmission
35:    Transmit analyzed results to the Cloud and output results to end-user device(s) with different visualizations.
36:  end while
37:end for
The AI-based real-time video solution seamlessly integrates with existing CCTV infrastructures, creating a Physical-Cyber-Physical (PCP) system that delivers actionable information to end users. Such a system consists of four key components: AI Modules, Server Module, Cloud Module, and End-user Devices, as illustrated in Fig. 2. Multiple camera streams can be fed into separate AI Modules in parallel. The object detector identifies bounding boxes for humans and other objects of interest, then forwards this bounding-box data to a tracker, which assigns unique IDs. Next, the tracked bounding boxes, including associated pixel data, are passed to a pose estimator to determine key skeletal joints. These skeletal data joints, organized into movement sequences, are subsequently processed by an anomaly detector to produce an anomaly score for the scene. All previously processed data are then sent to a feature-extraction stage to generate unique signatures for cross-camera identification, a task carried out by the server module. Finally, additional statistical data are analyzed and transmitted to the cloud module for end-user applications. The design of the AI and server modules was inspired by prior work [8], while the cloud module and user device components were adapted from another study [4]. Algorithm 1 provides more details regarding the complete SVS.
3.1 AI Module/Modules
@@
AI Module/Modules is a multi-stage pipeline optimized for real-time computer vision tasks. It processes image data in batches of 30 frames, in which the object detector identifies objects and the anomaly detector flags the behavior [8]. Object anomaly detection and behavioral anomaly detection [26, 27] are performed within this module respectively, with alerts sent to end-user devices for real-time response [12, 28].
One fundamental experiment was conducted to identify the optimal algorithm configuration for the AI node [8], specifically tailored to address real-world security requirements before going into further research. As highlighted by Pazho et al. [8], the GEPC method [29] is particularly effective for behavioral anomaly detection. ByteTrack [30] and OSNet [31] excel in pedestrian tracking and human re-identification, while the recent YOLOv8 [32] demonstrates superior performance in object detection and enhances human pose estimation through a top-down approach.
Such evaluation utilized the DukeMTMC dataset [33], with one-minute video clips of varying crowd densities. Initial validation was conducted on a laboratory server using default weights and parameters from original repositories, measuring throughput, latency, detection outputs, and computational resource utilization. The testing protocol employed two high-density scenarios, labeled ”Extreme” and ”Heavy,” each comprising 60-second videos at 60 frames per second, yielding 120 inference batches. Statistical validity was ensured by analyzing the middle 80 batches exclusively, eliminating potential bias from warm-up and cool-down periods in the performance metrics.
Four distinct methods underwent testing:
  • 1.
    YOLOv5+HRNet: The original system method
  • 2.
    YOLOv8pose: A top-down human pose estimation method.
  • 3.
    YOLOv8pose-p6: The most complex model of YOLOv8 for pose estimation.
  • 4.
    YOLOv8+HRNet: Object detection method with human pose estimation as a bottom-up method.
Table 1: Performance comparison among different system configurations. Data collected using lab server with a single node.
Crowd
Density
Method
Latency
FPS
Detections
GPU
Memory
CPU
Memory
Total
FLOPs
Extreme ($\sim$50 detects per second) YOLOv5+HRNet 14 14.82 1033 5.3 19.1 126.08
YOLOv8pose-p6 3.7 26.39 911 12.2 12.68 1067.38
YOLOv8pose 2.2 49.07 679 6.9 12.38 264.18
YOLOv8+HRNet 9.09 19 1217 5.5 19.3 182.18
Heavy ($\sim$20 detects per second) YOLOv5+HRNet 3.47 81.2 237 5.0 15.28 126.08
YOLOv8pose-p6 3.54 26.45 358 11.9 12.4 1067.38
YOLOv8pose 2.6 45.11 308 6.9 12.3 264.18
YOLOv8+HRNet 2.76 67.63 342 5.5 15.4 182.18
Table 1 evaluates system performance across varying crowd densities using four key metrics: average latency and throughput, total detection count, average GPU and CPU memory utilization, and the AI models’ total FLOPs. The combined YOLOv8+HRNet architecture demonstrated superior detection capabilities, particularly in challenging ”Extreme” and ”Heavy” crowd density scenarios, validating its robustness for real-world applications. For a comprehensive evaluation, our pipeline integrates YOLOv8 [32] for object detection, ByteTrack [30] for tracking, HRNet [34] for pose estimation, GEPC [35] for behavioral analysis, and OSNet [31] for re-identification.
Table 2: Accuray of different model weights implemented in the system
Model Dataset Metric Value
YoloV8l [32] COCO mAP 52.9
ByteTrack [30] MOT17 MOTA 80.3
HRNet_w32 [34] COCO AP 0.744
GEPC [29] NTU60 AUC-ROC 0.85
OSNet [31] OSNet R1 94.8
Importantly, all accuracy metrics in subsequent experiments are derived directly from the pre-trained model weights as published in each respective repository, without any additional fine-tuning or optimization. This approach ensures reproducibility and maintains consistency with the established benchmarks reported in the original publications. Table 2 summarizes the accuracy metrics of each model that is used for experiments on the existing well-established datasets.
3.2 Server Module